You have recently joined the finance department of Flames plc as a graduate trainee. The Finance Director would like to use the weighted average cost of capital as the discount rate to calculate whether to invest in a new overseas plant. She is comfortable calculating the cost of equity however she is unsure of how to calculate the cost of listed debt finance. You have been provided with the following information and asked to look into this.

The company has issued HK$50,000,000 of bonds (par value HK$1,000) which are redeemable at par in 3 years. These are currently trading at HK$ 1,015. The coupon payable is 8% annually and the tax rate is 25 %

The issued share capital of the company is HK$75,000,000 each with a nominal value of 50 cents. These are currently trading at 98 cents. The Finance Director has calculated the cost of equity to be 9%

  1. Discuss fully how the decision to go ahead with the new project would affect the market value of the company if the capital markets are;
    1. Strong form efficient
    2. Semi strong form efficient

Your answers should consider the assumptions underpinning the efficient market hypothesis
Answer #1

1) Weighted average cost of capital is nothing but the average of cost of debt and cost of equity by taking the market price as the weight of the respective costs .

Step 1: Cost of debt after considering tax benefit is : 8*0.75= 6%.

Market weight is HK$50,000,000/1000*1015= HK$50,750,000

Step 2: Cost of equity is 9%.

Market weight is HK$75,000,000/50*98=147,000,000.

Step 3: Weighted average of above two is:

(6*50,750,000+9*147,000,000)/197,750,000 =8.23%

2) The decision of the project depends on the NRV of the project by discounting it taking rate as 8.23%.

If the NRV is positive , company should go for the project or else not.

Strong form of efficient means the the market has the knowledge of even insider news and semi strong form of efficient means that the market doesn't have the insider information.

Answer: When markets are strongly efficient and the company goes for the project having positive NRV , the market value of the company will increase .

However , when market is semi strong efficient , there will not be any change in the market value of the company.